OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

OpenStudy (ajprincess):

Slope m of a line passing through two points \((x_1, y_1)\) and \((x_2,y_2)\) can be found using the formula \[m=\frac{y_2-y_1}{x_2-x_1}\] Does that help? @kevi23

OpenStudy (anonymous):

slo\[slope =\frac{ y }{ x}\] y2- y1/x2-x1

OpenStudy (anonymous):

A. -2 B. 1/2 C. -1/2 D.2

OpenStudy (anonymous):

a little @ajprincess

OpenStudy (anonymous):

Sure you can help a guy on slope but when it comes to pythareom theorem u cant help a person out man what a bunch of wussies

OpenStudy (anonymous):

there are the choices right there

OpenStudy (ajprincess):

\(x_1=6\),\(x_2=7\),\(y_1=2\) and \(y_2=4\)

OpenStudy (anonymous):

LOL!!!! @Rancd

OpenStudy (anonymous):

slope=m=y2-y1/x2-x1 Data: x1=6 x2=7 y1=2 y2=4 put the value in the formula m=7-6/4-2 m=-1/2 Option number C :)

OpenStudy (anonymous):

it cant be 4

OpenStudy (anonymous):

oh ok thank you @xoya, thank you to everyone who posted on here

OpenStudy (anonymous):

yeah c its easy wussies and u know the word im trying to say when it come to a hard problem im the besto

OpenStudy (anonymous):

yall were so helpfull

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yor welcome :) I hope you understood :)

OpenStudy (ajprincess):

Welcome:)

OpenStudy (anonymous):

Listen if u what to get hard problems got to this site brilliant.org to be the best schlorshipssss

OpenStudy (anonymous):

i do @xoya, but @Rancd u got me joked out, lmao ok at @Rancd

OpenStudy (anonymous):

what to become the best? try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

alright see u there at the top

OpenStudy (anonymous):

lol iight, ill see you there